Starting on January 1, 1987, and then on each January 1 until 1996 (10 payments), you will make payments of $1,000 into an investment which yields 10 percent. How much will your investment be worth on December 31 in the year 2006?   


1
Expert's answer
2021-04-08T07:18:10-0400



Balance by 31 Dec 2006

"A=P(1+ \frac{r}{n})" nt

"=16064.51(1+ \frac{0.1}{1})"10

"=41667.20169"